Find the potential inside and outside a spherical shell of radius R, which carries a uniform surface charge. Set the reference point at infinity


1
Expert's answer
2021-02-24T12:49:32-0500

The electric field inside:


"E=0"

The electric field outside:


"E=\frac{kq}{r^2},r>R"

"V(r)=-\int_\infty^rEdr=\frac{kq}{r},r>R\\V(r)=\frac{kq}{R},r<R"